A mixture of oxygen and methane gases contains oxygen at a partial pressure of 206 mm Hg and methane at a partial pressure of 668 mm Hg. What is the mole fraction of each gas in the mixture?

XO2 =
XCH4 =
Answer #1

Total pressure = 206+668

= 874 mm Hg

Xo2 = 206/874

= 0.236

Xco2 = 668/874

= 0.764